OpenStudy (anonymous):

sin2x-cos2x=(1+√3)/2

OpenStudy (tylerd):

do you mean \[\sin^2(x)-\cos^2(x)\] or \[\sin(2x)-\cos(2x)\]

OpenStudy (anonymous):

sin(2x)-cos(2x)

OpenStudy (tylerd):

ok one sec

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

\[\sin(2x)-\cos(2x)=\frac{(1+\sqrt3)}{2}\]Square both sides, then we have \[[\sin(2x)-\cos(2x)]^2=(\frac{1+\sqrt3}{2})^2\]Expand the left side, we have \[\sin^2(2x)-2\sin(2x)\cos(2x)+\cos^2(2x)=(\frac{1+\sqrt3}{2})^2\]Rearrange the terms, we have \[(\sin^2(2x)+\cos^2(2x))-2\sin(2x)\cos(2x)=(\frac{1+\sqrt3}{2})^2\]Use \(\sin^2a+\cos^2a=1\) and \(\sin(2a) = 2\sin a\cos a\) to simplify the left side, we have \[1-\sin(4x)=(\frac{1+\sqrt3}{2})^2\]Rearrange the terms again, we have \[\sin(4x)=1-(\frac{1+\sqrt3}{2})^2\] I guess you can do the rest :)

OpenStudy (anonymous):

Awesome!! thx a lot.. :)))

OpenStudy (anonymous):

wait a sec,what im doing now?i really hate this subject ;( can you tell me please?

OpenStudy (anonymous):

Which part do you not understand?

OpenStudy (anonymous):

sin(4x)=-0.866? if yes what now?

OpenStudy (tylerd):